# Resisting Contextual Interference in RAG via Parametric-Knowledge Reinforcement

### v3

@@ Appendix A Prompt designed in Knowledgeable-R1 and baselines
RAG prompting {search_results} You are a helpful assistant. After the user asks a question, you first think carefully and then give the answer. When responding, please keep the following points in mind: - The reasoning process and answer are enclosed within <think> </think> and <answer> </answer> tags, respectively. - Output your final answer directly between the tag <answer> </answer> without any intermediate steps. Here is an example: user’s question: what is the capital of China? <think> reasoning process here </think> <answer> BeiJing </answer>
@@
Appendix B Inference Input in Experiments
@@
Appendix C Experimental Setup
@@ Appendix D Evaluation
Appendix E Method Training Procedure
@@
$J_{\mathrm{RPK}}=\frac{1}{n_{pk}}\sum_{i=1}^{n_{pk}}\sum_{t=1}^{|o_{i}|}\min\left[\frac{\pi_{\theta}\left(o_{i,t}\mid p^{\prime},o_{i,<t}\right)}{\pi_{\theta_{\text{old}}}\left(o_{i,t}\mid p^{\prime},o_{i,<t}\right)}A_{i},\operatorname{clip}\left(\frac{\pi_{\theta}\left(o_{i,t}\mid p^{\prime},o_{i,<t}\right)}{\pi_{\theta_{\text{old}}}\left(o_{i,t}\mid p^{\prime},o_{i,<t}\right)};1-\epsilon,1+\epsilon\right)T(\hat{A}_{i})\right]$
@@
Appendix F Additional Experiment Results
@@
Method PC-MR PC-MC PC-QA NC-MR NC-MC NC-QA SC ExplainPE HotPotQA
2Wiki
MultiHopQA
Musique

### v4

@@ Appendix H Prompt designed in Knowledgeable-R1 and baselines
RAG prompting {search_results} You are a helpful assistant. After the user asks a question, you first think carefully and then give the answer. When responding, please keep the following points in mind: - The reasoning process and answer are enclosed within <think> </think> and <answer> </answer> tags, respectively. - Output your final answer directly between the tag <answer> </answer> without any intermediate steps. Here is an example: user’s question: what is the capital of China? <think> reasoning process here </think> <answer> BeiJing </answer> Question: {question}
@@
SFT baseline prompt You are a helpful assistant. After the user asks a question, you MUST directly give the answer. When responding, please keep the following points in mind: - Please answer with reference to the context, but if the context conflicts with your internal knowledge, prioritize your internal knowledge. - Output your final answer directly between the tag <answer> </answer> without any intermediate steps. Here is an example: Question: what is the capital of China? <answer> BeiJing </answer> Question: {question}
Appendix I Inference Input in Experiments
@@
Appendix J Experimental Setup
@@
Training stability. Knowledgeable-R1 remains stable across all our experiments: losses, rewards, and value estimates evolve smoothly across 200+ training runs covering both GRPO w/ RAG and Knowledgeable-R1 configurations. No collapse or divergence was observed, indicating that the additional PK/RPK objectives introduce no extra instability on top of the standard GRPO training loop.
Retrieval setup. We use dataset-provided contexts without any additional retriever. The contexts (including both correct and incorrect passages) are directly taken from the benchmark construction, without re-indexing or re-retrieval from a larger corpus. We adopt the top-5 paragraphs provided by these benchmarks, preserve their original order, and only truncate when the backbone model’s maximum length is exceeded.
Computational Efficiency. Knowledgeable-R1 uses 16 CK + 16 RPK trajectories (vs. GRPO’s 32). With shared forward passes, per-epoch time is 1h 38m vs. 1h 19m for GRPO — a modest $\sim$24% overhead for significantly larger robustness gains.
@@
Appendix L Method Training Procedure
@@
$J_{\mathrm{RPK}}=\frac{1}{n_{pk}}\sum_{i=1}^{n_{pk}}\sum_{t=1}^{|o_{i}|}\min\left[\frac{\pi_{\theta}\left(o_{i,t}\mid p^{\prime},o_{i,<t}\right)}{\pi_{\theta_{\text{old}}}\left(o_{i,t}\mid p^{\prime},o_{i,<t}\right)}T(\hat{A}_{i}),\operatorname{clip}\left(\frac{\pi_{\theta}\left(o_{i,t}\mid p^{\prime},o_{i,<t}\right)}{\pi_{\theta_{\text{old}}}\left(o_{i,t}\mid p^{\prime},o_{i,<t}\right)};1-\epsilon,1+\epsilon\right)T(\hat{A}_{i})\right]$
@@
Appendix M Additional Experiment Results
@@
Method PC-MR PC-MC PC-QA NC-MR NC-MC NC-QA SC ExplainPE HotPotQA 2Wiki MultiHopQA Musique